Solve the IVP -5 3 [= il (x1+[ 6te *) (O)=(-1) -3


a.
b.

Solve the IVP 79 [] = [-* ?] [?]+[6] [70] = [4] 1 0
Answer #1

[ ]]= [5 ] []+[t] [ce]-[ Let, A= Was find eigenvalue and eigenvector of A. 14-d1120 1-5-2 3 > (-5-2) (1-4) +9=0 +42 +5+=0 - 2Now solve for n, where n such that, [; })(09:11 -31; +3^2=1. ri ☆ +42 1-13 +22 ::7 22 choose, 22-01 we get n = : [1] = «[1]**..2A-30 =0 OF JE-887VE 2 ESO 2B = 6 > B = 3 2A-D=0 OFFE- st 166 ECO A =D. > A = -3 > D-3 -5C +3Fco -32+ f = 0. fc0 (01.- [***b) (%) =(-1] [2] = [:* ?][0][*] let, A= [31] Now find eigenvalue and eigenvector of A. I A-Il0 zali -3 (4-1)(1-2) +6=0 3 ² +3: []= [:]=»[] eigenvector corresponding eigenvalue dq=-1&s V22 2 ܙܝ - + - 19).=643*[] tee*[3] Let, At to ct to [3]. - ( 4 -]2 [2] [2] - eigenvector corresponding eigenualue dz=r2 is .. The general solution x-Itae [3] - set[] tie [1] + [3] YC) AC [1]